A U v ab x, x A V V 12 Real Money Income $800 XI $500 X2 $300-- 125 Leisure 80 90

a. In the above graph, if the price of consumption is $1, what is the wage rate the consumer is making while at X1?

b. In the above graph, how many hours is the individual working at x2?

c. Using the above information, does the income effect or substitution effect dominate moving from x1 to x2?
Answer #1

A) wage rate=(800-300)/125=500/125=4

Since 125 working hours can earn 500 which means wage rate is $4 per hour

b)At X2, leisure=90 and total number of hours=125. Thus number of hours individual work=125-90=35 hours

c) A decrease in wage rate will have two impact on leisure

opportunity cost of leisure decreases which increases consumption of leisure

decrease in wage rate will decrease real income and decrease the consumption of leisure

However net impact is an increase in leisure which means Substitution effect is greater than income effect.

substitution effect is dominating